OpenStudy (anonymous):

The value of 3pi/16 is a solution for 2cos^2(4x) -1 = 0?

OpenStudy (anonymous):

I know I have to change it to 2cos^2(4x) = 1

OpenStudy (anonymous):

but then I get stuck

OpenStudy (anonymous):

there is no need for this use double angle formula 2cos^2(x)-1 = cos2x

OpenStudy (anonymous):

it is given 2cos^2(4x) - 1= 0 cos(8x) = 0 do u follow the last step?

OpenStudy (anonymous):

yes I think so

OpenStudy (anonymous):

cos of 8 times 3pi/16?

OpenStudy (anonymous):

now cos8x = cos(npi/2) where n= odd numbers 8x = npi/2 in this case n=3 therefore 8x = 3pi/2 x = 3pi/16

OpenStudy (anonymous):

did u follow the steps?

OpenStudy (anonymous):

yes! thank you so much!! :)

OpenStudy (anonymous):

pleasure